# YUI: Day-ahead Electricity Price Forecasting Using Invariance Simplified Supply and Demand Curve

### v1

@@ 5.2. Forecasting the Supply Curve
We model daily supply curves using the most relevant historical data, discarding data prior to significant shape changes to ensure accuracy. When utilizing historical price equilibrium data to fit the supply curve prior to a specific target date, the volume of historical data naturally influences the quality of the fit. However, supply curves are more similar between closer dates. As depicted in Fig. 2(c), the degree of similarity between supply curves decreases when there is a substantial time gap between two dates. Consequently, our initial step is to identify and select the most suitable historical data for the fitting process. We model the supply curve for each day first, then examine the supply curves over these dates for significant shape changes. If the change surpasses a certain threshold, we only retain the data from the dates after the mutation.
@@
The necessity for the most recent dates when selecting historical data for fitting the supply curve limits the amount of usable data. So it is imperative to employ a model that is relatively simple and easy to fit when modeling the supply curve. We choose to represent the supply curve by a piecewise linear function, reflecting its phases that rise in different segments with different slopes. As previously mentioned, due to the operational characteristics of power plant generators (Roozbehani et al., 2022), the shape of the supply curve has consistently exhibited piecewise phases: rapid growth, steady growth, and then rapid growth again. Therefore, we employ the simplest form of a n-segment piecewise linear function to represent the supply curve. The slopes and intercepts of the n lines are represented by $w_{1}$, $w_{2}$, $\ldots$ , $w_{n}$ and $b_{1}$, $b_{2}$, $\ldots$, $b_{n}$ respectively. And $Q_{1}^{*}$, $\ldots$, $Q_{n-1}^{*}$ is the breakpoints of the piecewise linear function.
We optimize the parameters of this piecewise linear function using historical equilibrium points data. It’s recognized that numerous transaction time slots occur within a day. At each of these time slots, the day-ahead electricity price and total power generation across the market can be observed. These correspond to the scatter points on the supply and demand curve plane, note as $(Q_{1},P_{1})$, $(Q_{2},P_{2})$, $\ldots$, $(Q_{N},P_{N})$. The set of scatter points includes the closest $d$ days’ data, each day has $k$ time slots.
The goal is to fit these scatter points with the n-segment linear model such that the sum of the distances from all points to the n-segment line is minimized. The final constraint equation ensures the continuity of the supply curve.
(1) $\underset{n,w_{1},w_{2},\ldots,w_{n},b_{1},b_{2},\ldots,b_{n},Q_{1}^{*},\ldots,Q_{n-1}^{*}}{\text{minimize}}\sum_{j=1}^{N}d_{j}^{2}$
@@
$d_{j}=\left\{\begin{array}[]{ll}|P_{j}-(w_{1}Q_{j}+b_{1})|&\text{if }Q_{j}\leq Q_{1}^{*}\\ |P_{j}-(w_{2}Q_{j}+b_{2})|&\text{if }Q_{1}^{*}<Q_{j}\leq Q_{2}^{*}\\ \cdots&\cdots\\ |P_{j}-(w_{n}Q_{j}+b_{n})|&\text{if }Q_{j}>Q_{n-1}^{*}\\ j=1,...,N\\ \end{array}\right.$
$w_{1}*Q_{1}^{*}+b_{1}=w_{2}*Q_{1}^{*}+b_{2}$
@@
$w_{n-1}*Q_{n-1}^{*}+b_{n-1}=w_{n}*Q_{n-1}^{*}+b_{n}$
The issue of fitting piecewise linear lines has long been addressed with mature solutions. In this context, we refer to python package pwlf (Jekel and Venter, 2019), enabling us to swiftly approximate the supply curve. When fitting the supply curve on a specific dataset, sometimes the fitting is not accurate when using python package pwlf directly. In this case, we can add some small tricks to improve the accuracy of the fitting, the details are listed in Appendix B.
5.3. Forecasting the Demand Curve
The demand curve in the day-ahead electricity market, due to its price insensitivity, can be simplified as a price-axis parallel line, determined solely by the load rate at the target moment.
First, we approach the forecast of the target date’s available generation capacity from two angles: the correlation between capacity and other variables, and the temporal features inherent in the capacity itself. Given that the third party typically provides a capacity per day, the historical available generation capacity data is relatively small, rendering complex time-series forecast models unsuitable. So we opted for the XGBoost model (Chen and Guestrin, 2016) to establish the relationship between capacity and other variables. On trading day $D$, we have access to forecast values for weather conditions such as temperature and wind speed $\hat{Wd}_{h}^{D+x}$, as well as the forecast quantities of market demand $\hat{Qd}_{h}^{D+x}$. We train the XGBoost model using the historical data then employ the forecast values to forecast the target date’s capacity, note as $\hat{C}_{XGBoost}^{D+x}$. Additionally, we utilize TimesNet (Wu et al., 2022) to forecast the target date’s available generation capacity based on its own temporal features. TimesNet simplifies complex time series into periods, transforms them into two-dimensional space, and uses CNN for modeling. We use historical capacity data to forecast the target date’s capacity, note as $\hat{C}_{Timesnet}^{D+x}$.
Then, we integrate the two forecasts of the target date’s available generation capacity using a simple weighted average method. Weights are separately applied to the two forecasts, which are then summed to yield the final forecast, as illustrated in the following formula:
(2) $\hat{C}^{D+x}=\mu*\hat{C}_{Timesnet}^{D+x}+(1-\mu)*\hat{C}_{XGBoost}^{D+x}$
Subsequently, we calculate the forecast load rate by dividing the third-party forecast market demand quantities by the forecast available generation capacity. We then plot a line parallel to the price-axis on the supply-demand curve’s plane, where the load rate equals the forecast value. This approx the demand curve we aim to forecast.
6. EVALUATION
6.1. Datasets, Baselines and Experiment Settings
In our experiments, we utilized datasets from two distinct regions’ day-ahead electricity markets: Shanxi and ISO New England, some information is shown in Table.  1. The details are listed in the Appendix A.
Table 1. Shanxi and ISO New England Datasets Description.
@@ 6.1. Datasets, Baselines and Experiment Settings
Time span 2023/03/01 - 2023/10/31 2022/10/01 - 2023/09/30
Test 2023/04/01 - 2023/10/31 2023/01/01 - 2023/09/30
Forecsat target D+2 electricity price D+1 electricity price
Time slots per day 96 24
We selected two types of baselines: state-of-the-art models for time series forecasting and existing models for day-ahead electricity price forecasting. It worth noting that in order to simulate real market applications, our model adopts a daily rolling training method on the dataset, which means that new data is added every day for retraining. To align with our method, all baselines also adopt a daily rolling training method, updating the training set and validation set every day. Besides, a deep learning method uses a training set as long as YUI (only a few days) tends to yield subpar performance. In order to improve the accuracy of deep learning methods, we use as long a training set as possible during testing and adjust the hyperparameters when testing the baseline.
The state-of-the-art model in the current widely studied time series forecasting problems can be mainly divided into univariate forecasting and multivariate forecasting. We choose TimesNet (Wu et al., 2022), Koopa (Liu et al., 2023b) and Informer (Zhou et al., 2021) to represent the univariate forecasting models, and Autoformer (Wu et al., 2021), FEDformer (Zhou et al., 2022), DLinear (Zeng et al., 2023) and iTransformer (Liu et al., 2023a) to represent the multivariate forecasting models. The variables used in the multivariate forecasting models are exactly the same as those in the YUI, including historical price, capacity, demand data and forecast demand, weather data.
The existing engineering models focus on day-ahead electricity price forecasting can be divided into Statistical methods, Machine learning methods and Hybrid methods. Based on recommendations, we choose SARIMA (Zhao et al., 2017; McHugh et al., 2019) and Linear (Uniejewski et al., 2016; Lago et al., 2021) model to represent Statistical methods; Support Vector Machine (SVM) (Che and Wang, 2010; Wang et al., 2017; Prahara and Hariadi, 2022), XGBoost (Manfre Jaimes et al., 2023; Xie et al., 2022b) and Deep Neural Network (DNN) (Yamin et al., 2004; Lago et al., 2018a; Darudi et al., 2015) to represent Machine learning methods; and LASSO-RF (Ludwig et al., 2015), VMD-LSTM (Xiong and Qing, 2023) to represent Hybrid methods. More baseline details are in Appendix C.
@@
6.2. Evaluation Metric
In the field of electricity price forecasting, the most widely used metrics to measure the accuracy of point forecasts are the Mean Absolute Error (MAE), the Root Mean Square Error (RMSE), and the Mean Absolute Percentage Error (MAPE). In particular, in most electricity trade applications, the underlying risk, profits, and costs depend linearly on the price and on the forecasting errors. Hence, linear metrics represent better than quadratic metrics the underlying risks of forecasting errors (Lago et al., 2021).
The formula for sMAPE is:
(3) $sMAPE=\frac{100\%}{n}\sum_{t=1}^{n}2\frac{|Y_{t}-\hat{Y}_{t}|}{|Y_{t}|+|\hat{Y}_{t}|}$
6.3. Results and Analysis
Table 2. Results on Shanxi and ISO New England Datasets.
@@ 6.3. Results and Analysis
MAE(¥/MWh) sMAPE MAE(\$/MWh) sMAPE
TimesNet 69.351 0.30108 13.787 0.31880
Koopa 74.500 0.28235 11.316 0.25691
Informer 103.34 0.33552 11.918 0.29089
Autoformer 85.920 0.30800 17.548 0.45804
iTransformer 74.983 0.29092 11.147 0.27245
DLinear 69.096 0.27178 13.315 0.31507
FEDformer 80.157 0.31438 16.740 0.39352
SARIMA 70.777 0.27347 9.3656 0.21114
Linear 85.450 0.29371 24.584 0.57063
SVM 86.915 0.30923 16.257 0.36596
XGBoost 63.350 0.23115 10.727 0.23246
DNN 98.922 0.33782 18.403 0.36594
Lasso-RF 62.198 0.24801 9.3005 0.20008
VMD-LSTM 59.245 0.23827 23.536 0.53750
YUI 51.065 0.16481 7.8839 0.17303
YUI-XGB 55.283 0.17601 11.976 0.23657
YUI-MLP 60.837 0.18008 9.0509 0.21258
YUI-PLOY 53.779 0.16339 10.602 0.21690
YUI-EXP 53.106 0.17138 10.698 0.20753
6.3.1. Main results
In Table. 2, we report the MAE and sMAPE on the Shanxi dataset since YUI model is deployed, as well as the results of MAE and sMAPE on a public dataset, ISO New England. Table. 2 shows that YUI’s MAE is lower than all other methods and YUI’s sMAPE is the lowest among all the methods. This indicates that our method outperforms all other methods in terms of forecast accuracy. It worth noting that even when compared to the second-best method, our method still manages to reduce the MAE by approximately $13.81\%$. On the sMAPE metric, the performance of YUI surpasses the second by $28.7\%$.
(a) When a generic time series forecasting model is directly applied to the task of forecasting day-ahead electricity prices, it often encounters difficulties in accurately forecasting high prices.
(b) Models applied in the industry, demonstrate relative success in forecasting trends of price increases, although the extent of these increases is somewhat limited.
(c) The YUI model has lowest MAE during these high-price time slots.
Figure 5. Day-ahead electricity prices and forecast values for the 24 time slots of ISO New England on July 17, 2023 and MAE for high-price time slots.
To help reproduce our results, we also experiment on an open dataset. The results on the ISO New England dataset are similar, showing strong robustness. YUI continues to excel by achieving the lowest forecast error once again, demonstrating a reduction in forecast error by $15.2\%$ in terms of MAE and $13.5\%$ in sMAPE.
6.3.2. Analysis of capacity forecast accuracy
Table 3. Capacity Forecast Results.
YUI in Shanxi YUI in ISO Official Forecast in ISO
RMSE 1556.4 863.34 1899.0
MAE 1200.2 675.26 1713.8
sMAPE 0.032269 0.033450 0.091486
When we shift the supply-demand curve’s horizontal coordinate from quantity to load rate, we forecast a new variable: available generation capacity. There’s a concern that this new data might add to cumulative forecast errors. We can initially address this skepticism through our final price forecasts’ accuracy. We can also validate our method by checking our capacity forecast’s precision. The ISO New England dataset conveniently includes the ISO organization’s capacity forecasts, allowing us to compare directly with our forecast results. The results from our models, as shown in the Table. 3, demonstrate that our model exhibits exceptional accuracy in capacity forecasting, achieving a mere $3\%$ sMAPE error, thereby introducing minimal error. Furthermore, when compared horizontally, our model significantly outperforms the official forecast provided by ISO New England. Therefore, we believe that our method of forecasting the available generation capacity is both valid and highly effective.
6.3.3. Analysis of forecast reliability
In order to ensure the reliability of the forecast, we need to verify that the method maintains high forecast accuracy at different times. We divide the test set of the ISO dataset (total length of 9 months) independently by month, and calculate the forecast error (MAE, sMAPE) of each method separately each month. A good robust method should maintain similar forecast accuracy in each month. We collect the forecast errors of different months, calculate the variance of the errors of each method, and the results show that YUI is the most stable in prediction accuracy. Please refer to Tab. 4.
Table 4. Variance of Forecast Errors of Different Months.
ISO New Engalnd
VAR4MAE VAR4sMAPE
Linear 174.079452 0.006546
XGBoost 42.249195 0.005145
DNN 217.487628 0.021724
SARIMA 38.154287 0.00304
SVM 101.904932 0.012685
Koopa 118.622141 0.008756
iTransformer 119.223867 0.011642
DLinear 98.435843 0.006762
TimesNet 259.261455 0.025859
VMD-LSTM 78.225068 0.014499
FEDformer 202.407886 0.010989
Lasso-RF 39.060255 0.003403
YUI 38.316854 0.002448
6.3.4. Ablation study
@@
In the part of forecasting the supply curve, we employ a piecewise linear function with n segments to fit the supply curve. This approach allows us to accurately represent the different phases of the supply curve, each characterized by varying slopes. The supply curve can also be fitted in other ways, such as cubic functions, exponential functions, or even using models like MLP, XGBoost to implicitly learn the relationship between price and load rate. To validate the appropriateness of our fitted function, we conducted tests on several variants of the YUI model: 1) YUI-XGB: using the XGBoost model to implicitly model the supply curve; 2) YUI-MLP: using the MLP model; 3) YUI-PLOY: using a cubic function; 4) YUI-EXP: using an exponential function. Apart from the different supply curve fitting functions used, these models have no other differences from YUI. The results of our ablation study, as shown in the Table. 2, indicate that the original YUI model outperforms all variants on both datasets. This outcome validates the effectiveness of our design choice to use a piecewise linear function for fitting the supply curve.
6.4. Case Study
We utilize publicly available dataset ISO New England to scrutinize the specific conditions under which our model excels. The distribution of day-ahead electricity prices in the dataset is somewhat dispersed in the high-price range, with 75% of prices falling below \$40/MWh, while the peak price can reach up to \$300/MWh. This type of information is infrequent and lacks a clear time series pattern, leading to subpar performance of baseline models, particularly time series forecasting models, in the high-price range (Rahaman et al., 2019; Xu, 2020). We have visualized the data from the ISO dataset dated July 17, 2023. Fig. 5(a) illustrates that a generic time series forecasting model, when directly applied to the problem of day-ahead electricity prices forecasting, struggles to accurately forecast high prices. Conversely, industry-applied models, as depicted in Fig. 5(b), are relatively more successful in forecasting price increase trends, albeit with limited magnitude. The YUI model effectively translates the electricity price from the time domain to the supply domain. Infrequent high prices in the time domain, which are less dispersed, are represented on the supply curve as linear segments with fewer but regularly shaped sample points. Consequently, the forecasts align more closely with the actual electricity prices. Furthermore, we compile all the high-price periods in the test set where the real day-ahead electricity price exceed \$40/MWh. The YUI model outperforms other models during these high-price periods, as demonstrated in Fig. 5(c).
6.5. Possible Reasons for Good Performance
YUI, which forecasts day-ahead electricity prices using a simplified supply and demand curve, has demonstrated impressive performance. We attribute this success to several key factors:
Incorporation of Prior Knowledge. Traditional data-driven models might find it hard to grasp the economic rules of supply and demand that guide price trends. YUI, on the other hand, is naturally split into two parts: one for forecasting the supply curve and another for the demand curve. This setup helps us more effectively use the processes of price formation.
Simple Model with Shorter Training Sets to Mitigate Data Drift. As depicted in Fig. 2(c), the similarity between supply curves diminishes when the time interval between two dates extends. To model the supply curve with precision, we utilize a piecewise linear function. This streamlined model negates the necessity for a large training set, thereby eliminating potential discrepancies between the fitted pattern and the target date. YUI employs historical data from merely a week prior to the target date for modeling the supply curve. In contrast, data-driven models that necessitate longer training sets, typically encompassing several months of data, are unable to circumvent the phenomenon of data drift.
Problem Decomposition. We simplify a complex time-series forecasting problem into an easier one, using a forecast for a more predictable variable and a basic univariate regression. In the electricity market, forecasting demand is simpler than price due to its stability. Electricity prices, influenced by various factors, are more volatile, making forecasting challenging. The detailed comparison result of predictability can be seen in Appendix D. Our approach is similar to methods like Autoformer (Wu et al., 2021) and FEDformer (Zhou et al., 2022), which decompose time series into components with obvious seasonality and trend, forecast them separately, and combine the results. These methods transition the price from the time domain into the frequency domain, while YUI offers a new perspective on this by focusing on the supply-demand relationship.
7. CONCLUSION
This paper introduces an innovative method for forecasting prices in the day-ahead electricity market. By employing the principles of time invariance and price insensitivity, we streamline the forecast of supply and demand curves. In scenarios where traditional economic modeling methods falter due to data scarcity, YUI effectively forecasts prices utilizing simplified supply and demand curves. Rigorous testing on two electricity datasets reveals that our model surpasses existing top-tier methods, highlighting the potential of YUI to enhance reliability in the day-ahead electricity market operations. Future endeavors will concentrate on refining our model and investigating its relevance to other markets. This research signifies a crucial stride towards a more sustainable and efficient energy future.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
Appendix A Datasets Introduction
In these two datasets, we use the supply and demand quantities data of thermal power generation to construct the supply and demand curve. The day-ahead electricity market comprises several key supplier categories: thermal, wind, solar, nuclear, and hydro power generation (Bichler et al., 2022). The supply curves for these categories exhibit significant variations. The day-ahead electricity market can be approximated as a perfectly competitive market, (OpenStax, 2024d, c, b, a) where highest marginal costs equals the price (Hogan, 2022). In the electricity market, the highest marginal cost is typically associated with thermal power generation. (Su et al., 2020; Hogan, 2022) The ISO New England dataset corroborates this perspective. It documents the supply costs of various energy sources across 108,646 trading intervals from October 1, 2022, to November 11, 2023. The data reveals that in 80.1% of these intervals, the marginal cost is tied to thermal power generation. This suggests a significant role of thermal power in influencing market dynamics.  (iso, 2023)This observation aligns well with conventional wisdom for several reasons:
  • Historical precedence. Thermal power generation was established earlier than other forms of energy generation, providing a stable power supply (Yang and Xu, 2022; Farnoosh, 2022).
  • Market dominance. Thermal power holds a substantial share of the electricity market, reflecting the industry’s maturity (Liu et al., 2021).
  • Resource allocation. The cost of thermal power generation, which relies on higher-cost fossil fuels, exceeds that of photovoltaic, solar, and hydro power generation (Liu et al., 2020).

### v2

@@ Introduction
We use the most relevant historical data to ensure accuracy. Supply curves are more similar between closer dates, as shown in Figure 4. Therefore, our initial step is to identify and select the most suitable historical data. We model the supply curve for each day and examine these curves for deviations in shape. If a deviation exceeds a threshold, we retain only the data after the deviation.
@@
The necessity for the most recent dates limits the amount of usable data, so it is imperative to employ a model that is relatively simple and easy to fit the supply curve. We represent the supply curve with a piecewise linear function, reflecting its phases with different slopes. Due to power plant generator characteristics (Roozbehani et al. 2022), the supply curve consistently exhibited piecewise phases: rapid growth, steady growth, and then rapid growth again. Therefore, we employ the simplest form of a n-segment piecewise linear function to represent the supply curve, named CoPiLinear. The slopes and intercepts of the n lines are represented by $w_{1}$, $w_{2}$, $\ldots$ , $w_{n}$ and $b_{1}$, $b_{2}$, $\ldots$, $b_{n}$ respectively. And $LR_{1}^{*}$, $\ldots$, $LR_{n-1}^{*}$ is the breakpoints of the piecewise linear function.
We optimize the parameters using historical equilibrium points data. Throughout the day, we can observe prices and load rates at various transaction time slots. These correspond to scatter points on the supply and demand curve plane, note as $(LR_{1},P_{1})$, $(LR_{2},P_{2})$, $\ldots$, $(LR_{N},P_{N})$. The set of scatter points includes the closest $d$ days’ data, each day has $k$ time slots. The goal is to fit these scatter points with CoPiLinear, minimizing the sum of distances from all points to the n-segment line. The final constraint equation ensures the continuity of the supply curve.
$\underset{n,w_{1},w_{2},\ldots,w_{n},b_{1},b_{2},\ldots,b_{n},LR_{1}^{*},\ldots,LR_{n-1}^{*}}{\text{minimize}}\sum_{j=1}^{N}d_{j}^{2}$ (1)
@@
$d_{j}=\left\{\begin{array}[]{ll}|P_{j}-(w_{1}LR_{j}+b_{1})|&\text{if }LR_{j}\leq LR_{1}^{*}\\ |P_{j}-(w_{2}LR_{j}+b_{2})|&\text{if }LR_{1}^{*}<LR_{j}\leq LR_{2}^{*}\\ \cdots&\cdots\\ |P_{j}-(w_{n}LR_{j}+b_{n})|&\text{if }LR_{j}>LR_{n-1}^{*}\\ j=1,...,N\\ \end{array}\right.$
$w_{1}*LR_{1}^{*}+b_{1}=w_{2}*LR_{1}^{*}+b_{2}$
@@
$w_{n-1}*LR_{n-1}^{*}+b_{n-1}=w_{n}*LR_{n-1}^{*}+b_{n}$
Fitting piecewise linear lines is a well-established problem with mature solutions. We use the Python package pwlf (Jekel and Venter 2019) to quickly approximate the supply curve. However, direct application that package on certain datasets may result in inaccuracies. To enhance fitting accuracy, we employ specific techniques detailed in Appendix C.
Forecast related variables
Electricity demand fluctuates significantly over time, making the demand curve challenging to forecast. By forecasting the supply curve, we can determine the price by identifying the demand quantity at the intersection of the supply and demand curves, without needing to forecast the entire demand curve.
Next, we need to forecast key supply-demand variables. The forecasted capacity data is used to adjust the horizontal axis of the CoPiLinear model from load rate to supply volume, creating the actual supply curve. The forecasted demand volume is then input into the CoPiLinear model to obtain the forecasted price.
These variables exhibit strong periodicity, making TSF models particularly effective, as detailed in Appendix D. Given their importance for the market, numerous mature TSF methods exist. Many service providers and market operators offer forecast data to aid market participants’ decision-making, so we directly utilize publicly available forecasts from third parties. In markets lacking forecasts for certain variable, such as capacity, we employ a simple model, Random Forest, to generate the necessary forecasts.
Evaluation
Datasets and Baselines
In our experiments, we utilized datasets from two distinct regions’ day-ahead electricity markets: Shanxi and ISO New England, some information is shown in Table 2. The details are listed in the Appendix E.
@@
Time span 2023/03/01 - 2024/3/31 2022/10/01 - 2023/12/31
Test 2023/04/01 - 2024/3/31 2023/01/01 - 2023/12/31
Time granularity 15min 1hour
Table 2: Shanxi and ISO New England Datasets Description.
We selected two types of baselines: models relying solely on temporal dependence and those capturing variable correlations. For temporal dependence, we chose domain-specific methods such as SARIMA (Zhao et al. 2017) and VMD-LSTM (Xiong and Qing 2023), and generic TSF methods like DLinear (Zeng et al. 2023) and PatchTST (Nie et al. 2022). For variable correlations, we selected domain-specific methods including Linear (Uniejewski, Nowotarski, and Weron 2016), SVM (Prahara and Hariadi 2022), XGBoost (Manfre Jaimes et al. 2023), DNN (Lago, De Ridder, and De Schutter 2018), and Lasso-RF (Ludwig, Feuerriegel, and Neumann 2015), along with generic TSF methods such as Koopa (Liu et al. 2023b), TSMixer (Chen et al. 2023b), Informer (Zhou et al. 2021), Autoformer (Wu et al. 2021), FEDformer (Zhou et al. 2022), Crossformer (Zhang and Yan 2023), and iTransformer (Liu et al. 2023a). The variables used in these generic TSF models are exactly the same as those in the CoPiLinear, including historical price, capacity, demand data and forecast demand, weather data.
To simulate real market applications, our model uses a daily rolling training method, adding new data each day for retraining. All baselines also adopt this method, updating the training and validation sets daily. Deep learning methods with short training sets tend to perform poorly. To improve accuracy, we use the longest possible training set and adjust hyperparameters during testing. More baseline details are in Appendix F.
Experiment Settings and Evaluation Metric
@@
Based on related work (Lago et al. 2021), we select MAE and sMAPE to evaluate forecast accuracy. Reference (Zheng et al. 2020) guides our choice of $R$ and $\alpha$ to assess the revenue of the forecast. Detailed descriptions of the evaluation metrics are provided in Appendix G.
Results and Analysis
@@
MAE(\$/MWh) sMAPE Revenue(\$/year) $\alpha$ $MAE_{CS}$ MAE(\$/MWh) sMAPE Revenue(\$/year) $\alpha$ $MAE_{CS}$
SARIMA 13.0 0.383 8.40 ×$\times 10^{7}$ 0.436 16.1 8.97 0.211 3.43 ×$\times 10^{7}$ 0.447 13.0
VMD-LSTM 12.7 0.357 8.48 ×$\times 10^{7}$ 0.440 16.7 21.7 0.510 1.94 ×$\times 10^{7}$ 0.227 24.6
DLinear 12.6 0.370 8.40 ×$\times 10^{7}$ 0.443 15.1 12.0 0.292 2.97 ×$\times 10^{7}$ 0.368 17.9
PatchTST 13.0 0.388 8.32 ×$\times 10^{7}$ 0.437 16.3 9.75 0.225 3.27 ×$\times 10^{7}$ 0.423 17.2
Linear 18.0 0.440 7.25 ×$\times 10^{7}$ 0.367 17.4 31.7 0.669 1.27 ×$\times 10^{7}$ 0.163 39.7
SVM 12.2 0.349 8.59 ×$\times 10^{7}$ 0.448 15.6 14.6 0.339 2.66 ×$\times 10^{7}$ 0.316 19.5
XGBoost 11.2 0.317 8.76 ×$\times 10^{7}$ 0.466 12.4 9.57 0.219 3.32 ×$\times 10^{7}$ 0.430 14.5
DNN 14.4 0.409 7.98 ×$\times 10^{7}$ 0.415 15.8 15.8 0.327 2.40 ×$\times 10^{7}$ 0.295 20.1
Lasso-RF 11.4 0.355 8.76 ×$\times 10^{7}$ 0.462 12.7 8.80 0.196 3.52 ×$\times 10^{7}$ 0.454 11.2
Koopa 13.2 0.387 8.24 ×$\times 10^{7}$ 0.433 16.6 10.5 0.246 3.26 ×$\times 10^{7}$ 0.406 16.2
TSMixer 12.4 0.384 8.51 ×$\times 10^{7}$ 0.435 15.8 12.8 0.306 2.86 ×$\times 10^{7}$ 0.349 19.5
Informer 16.6 0.414 7.43 ×$\times 10^{7}$ 0.384 18.7 11.4 0.285 3.00 ×$\times 10^{7}$ 0.381 17.1
Autoformer 16.1 0.431 7.62 ×$\times 10^{7}$ 0.391 22.7 15.8 0.416 2.44 ×$\times 10^{7}$ 0.297 24.4
FEDformer 14.2 0.411 8.40 ×$\times 10^{7}$ 0.419 18.1 14.9 0.377 2.64 ×$\times 10^{7}$ 0.307 26.2
Crossformer 11.5 0.361 8.36 ×$\times 10^{7}$ 0.462 12.7 10.2 0.238 3.14 ×$\times 10^{7}$ 0.412 15.0
iTransformer 13.1 0.394 8.32 ×$\times 10^{7}$ 0.436 15.7 10.3 0.258 3.29 ×$\times 10^{7}$ 0.408 17.7
CoPiLinear 8.61 0.196 9.49 ×$\times 10^{7}$ 0.512 10.8 7.77 0.175 3.62 ×$\times 10^{7}$ 0.491 9.78
Co-POLY 9.30 0.203 9.30 ×$\times 10^{7}$ 0.500 12.0 10.6 0.222 3.17 ×$\times 10^{7}$ 0.400 15.2
Co-EXP 9.68 0.227 9.30 ×$\times 10^{7}$ 0.494 11.5 10.6 0.213 3.06 ×$\times 10^{7}$ 0.398 15.7
Co-XGB 10.1 0.245 9.20 ×$\times 10^{7}$ 0.484 13.1 12.1 0.241 2.85 ×$\times 10^{7}$ 0.361 15.8
Co-MLP 11.8 0.284 8.68 ×$\times 10^{7}$ 0.457 13.1 12.7 0.315 2.82 ×$\times 10^{7}$ 0.352 16.4
Table 3: Results on Shanxi and New England Datasets.
Main results
In Table 3, we present the main results. For consistency, prices in the Shanxi dataset are converted to US dollars using the exchange rate of 7.16 RMB to 1 USD as of the writing date. For the Shanxi dataset, CoPiLinear achieves the lowest MAE and sMAPE, outperforming the second-best by 22.9% in MAE and 38.0% in sMAPE. This superior forecast accuracy also boosts electricity purchase revenue, with CoPiLinear generating the highest revenue and alpha. To ensure reproducibility, we also test on the ISO New England dataset, where CoPiLinear again shows the lowest forecast error and achieving the highest revenue. Note that the electricity volume consumed by residents in Shanxi is approximately 3 times that of ISO New England, resulting in higher revenue.
Analysis of forecast accuracy continuity
Figure 6: Shanxi market transaction process.
To ensure forecast reliability, we verify that the method maintains high accuracy across different times. We divide the ISO dataset’s test set (1 year) by month, select the top five methods with the best overall performance, and calculate each method’s monthly forecast error (MAE). A robust method should exhibit consistent accuracy each month. We plot the monthly forecast errors as box plots in Figure 6, revealing that CoPiLinear is the most stable. Additionally, higher MAEs are observed in summer months, coinciding with increased market energy consumption and price.
Ablation study 1
@@
For fitting the supply curve, we use a piecewise linear function, accurately capturing the different phases characterized by varying slopes. Alternative fitting methods include cubic functions, exponential functions, or models like MLP and XGBoost to implicitly learn the price-load correlation. To validate our approach, we tested several CoPiLinear variants: 1) Co-PLOY: cubic function; 2) Co-EXP: exponential function; 3) Co-XGB: XGBoost model; 4) Co-MLP: MLP model. These models differ only in the supply curve fitting function. Our ablation study results Table 3 show that the original CoPiLinear model outperforms all variants on both datasets, confirming the effectiveness of using a piecewise linear function.
Ablation study 2
SARIMA TimesNet XGBoost Random Forest Official Forecast
RMSE 2030.4 1532.2 889.07 839.34 1899.0
MAE 1450.8 1182.3 698.44 651.41 1713.8
sMAPE 0.070585 0.57557 0.035612 0.33041 0.091486
Table 4: Capacity Forecast Results.
When shifting the supply curve’s horizontal axis from quantity to load rate, we use a Random Forest model to forecast capacity. Given the limited capacity data (one entry per day), we select models with smaller training set requirements to compare forecasting accuracy: SARIMA, TimesNet and XGBoost. What’s more, the ISO New England dataset includes the organization’s official capacity forecasts, allowing direct comparison. As shown in Table 4, our model achieves exceptional accuracy with only a 3% sMAPE error, significantly outperforming the official forecast. Our Random Forest model’s higher accuracy compared to other models further validates the design of this module. This demonstrates the validity and effectiveness of our capacity forecasting method.
Case Study
Figure 7: Shanxi market transaction process.
Our method outperforms existing approaches in capturing the correlation between supply-demand and price. This allows for more sensitive adjustments to price forecasts based on forecasted demand or other variables, especially during significant supply and demand fluctuations. Figure 3 analyzes the challenging high prices in the ISO New England market on July 17, 2023, due to a sharp increase in forecasted demand. Figure 7 shows the CoPiLinear forecast results align more closely with actual prices compared to other methods. The CoPiLinear forecasted price peaks a few phases earlier than the actual prices, it can be explained that it synchronizes completely with the demand surge, whereas the energy storage system in reality delays the impact of the demand surge on prices.
We validate this in two datasets. If the forecasted demand during the day exceeds the previous week’s maximum or falls below the minimum, we anticipate significant price fluctuations. We selected these dates and compared the MAE of various forecasting methods, as shown in the $MAE_{CS}$ column of Table 3. All methods showed increased MAE on these dates, with CoPiLinear achieving the highest accuracy in both datasets.
CONCLUSION
This paper has highlighted the limitations of current day-ahead electricity price forecasting methods and introduced a novel approach to fit supply curve. By assuming short-term stability, CoPiLinear has significantly improved accuracy, benefiting residents. Rigorous testing on two electricity datasets shows that CoPiLinear outperforms top-tier methods, underscoring its potential to reduce residents’ costs and enhance forecast reliability in day-ahead market operations. Future work will refine the model and explore its applicability to other markets. This achievement offers new perspectives for electricity price forecasting research and showcases the potential of artificial intelligence in the socio-economic field.
Appendix A
High prices are mainly driven by high demand, which can help forecast these irregularities. However, TSF models struggle to capture the correlation between prices and supply-demand dynamics. According to the price formation mechanism, forecasted demand should significantly impacts forecasted price. Yet, even with substantial changes in demands, existing models show minimal price variation. This inconsistency with real market behavior indicates that these models do not adequately incorporate supply and demand variables, leading to poor forecasting performance. See Appendix A for detailed experiment results.
To investigate this, we alter the demand for the target date, setting it to zero or doubling it, while keeping other inputs constant. Using data from Shanxi province for January 2024, we examine the impact of significant demand changes on the forecasted prices of four models: DNN, Lasso-RF, iTransformer, and Crossformer. The first two models are commonly used in day-ahead electricity price forecasting, while the latter two are state-of-the-art (SOTA) TSF models that consider relationships between variables. The sMAPE between the forecasted prices before and after the change is generally small, not exceeding 20%, as shown in Table. 5. Figure 8 shows the forecasted prices for January 15th, with minimal changes even when demand is significantly altered. Interestingly, forecasted prices often increase when demand is zero and decrease when demand is doubled, contradicting economic theory.
DNN Lasso-RF iTransformer Crossformer
Zero Demand 0.20077 0.10006 0.055960 0.13900
Double Demand 0.27866 0.15165 0.073360 0.078306
Table 5: sMAPE between the forecasted prices before and after demand change.
(a) Domain-specific models.
(b) SOTA TSF models.
Figure 8: Forecasted prices with altered demand for January 15th, 2024.
Appendix B